solve this pair of equations by substitution
4x+y=-19 7x-y=8

Answer by rapaljer(4551)

To solve by substitution, you have to solve for one of the variables in terms of the other. Probably the first equation is the easiest to solve for y.

4x+y=-19
y = -4x - 19

Substitute this into the second equation for y:
7x-y=8
7x - (______) = 8
7x - (-4x - 19) = 8
7x + 4x + 19 = 8
11x + 19 = 8
11x + 19 - 19 = 8-19
11x = -11
x= -1

Now, substitute this value of x into the y = equation to solve for y:
y = -4x - 19
y = -4(-1) - 19
y = 4 -19
y=-15

Check by substituting into the OTHER equation:
7x - y = 8
7(-1)-(-15) = 8
-7+15 = 8
8=8
It checks!

R^2 at SCC